OpenStudy (erinweeks):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n2 - n + 2

OpenStudy (anonymous):

Start by proving the base case: \(n=1\)

OpenStudy (erinweeks):

i dont know how too.

OpenStudy (anonymous):

\[ (1)^2-(1)+2 \]

OpenStudy (anonymous):

Show that it is a factor of \(2\)

OpenStudy (erinweeks):

of so'd it be (2)^2 - (2) + 2 ?

OpenStudy (erinweeks):

4 - 2 +2 = 4?

OpenStudy (anonymous):

No, show that when \(n=1\) it's a factor of \(2\).

OpenStudy (erinweeks):

how do i that?

OpenStudy (anonymous):

We are beyond high school now. We know how to substitute values into variables.

OpenStudy (erinweeks):

so it'd be 2 - 1 + 2? = 3

OpenStudy (anonymous):

Okay so we have:\[ n^2 - n + 2 \]What does it equal when \(n=1\)?

OpenStudy (erinweeks):

1^2 - 1 + 2 = 2 - 1 +2 = 3?

OpenStudy (anonymous):

\(1^2\neq 2\)

OpenStudy (erinweeks):

okay so the statement isnt true?

OpenStudy (anonymous):

The problem is that you aren't doing the exponent correctly.\[ 1^2=1\times 1=1 \]

OpenStudy (anonymous):

Wait, are you trying to make fun of me?